OpenStudy (anonymous):

help please Add the fractions. Select the answer that is in simplest form. 2/5 + 3/10=

OpenStudy (anonymous):

get the fractions the the same denominator, add then simplify

OpenStudy (anonymous):

im lost

OpenStudy (anonymous):

so 2/5=4/10 4/10+3/10=7/10

OpenStudy (anonymous):

7/10

OpenStudy (anonymous):

kan u explain how u did that?

OpenStudy (anonymous):